From a group of 7 men and 5 women a 4 person committee is chosen what is the probability that all 4 are men or all 4 are women?


1
Expert's answer
2021-05-02T15:44:34-0400

The number of all possible events is a number of ways we can choose 4 people out of 12 which is

"C(12,4) = \begin{pmatrix}
 12 \\ 4
\end{pmatrix} = 495"

The number of ways we can choose all women in the committee

"C(5,4) = \begin{pmatrix}
 5 \\ 4
\end{pmatrix} = 5"

The number of ways we can choose all men in the committee

"C(7,4) = \begin{pmatrix}
 5 \\ 4
\end{pmatrix} = 35"

So the probability is

"\displaystyle P(A) = \frac{5}{495} + \frac{35}{495} = \frac{40}{495}=0.08 =8\%"


Answer: 8%